An object has a volume of 4.0 m3 and weighs 40,000 N. What will its weight be in water? Select one: a. 40,000 N b. 39,200 N c. 9,800 N d. 800 N
Answer #1

Weight of the object in water = Weight in air - buoyant Force

According to Archimedes Principle, Buoyant force is equal to weight of the water displaced.

W_{water}= W_{air}-V\rho _{w}g  

W_{water}= 40,000 N-4\times 1000\times 9.8 N = 800 N